a standard deck of cards contains 52 cards. one card is selected from the deck.

a) compute the probability of randomly selecting a ten or three

b) compute the probability of randomly selecting a ten or three or seven

c) Computer probability of randomly selecting a king or spade


1
Expert's answer
2021-09-21T00:29:09-0400

a) "P=\frac{4+4}{52}=\frac{2}{13}\approx 0.1538."


b) "P=\frac{4+4+4}{52}=\frac{3}{13}\approx 0.2308."


c) "P=\frac{4+13-1}{52}=\frac{4}{13}\approx 0.3077."